Dallas Engineering purchased a machine five years ago at a cost of $484,000. The machine is being depreciated using the straight-line method over eight years. The tax rate is 25 percent and the discount rate is 13 percent. If the machine is sold today for $209,000, what will the aftertax salvage value be?

$202,125

$227,485

$214,500

$194,000

$186,775
Answer #1

$202,125

After tax salvage value = Salvage value-((Salvage value - Book Value as of today)*Tax Rate)
= 209000-((209000-181500)*25%)
= $       2,02,125
Working:
# 1
Straight line depreciation = Cost / Useful Life
= 484000/8
= $           60,500
# 2
Cost = $       4,84,000
Less Accumulated depreciation = $           60,500 * 5 = $       3,02,500
Book Value as of today = $       1,81,500